find the equation of the line that passes through the point (2,9) and has slope 4. write your answer in point-slope form.
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20062)

find the equation of the line that passes through the point (2,9) and has slope 4. write your answer in point-slope form.
Substitute (x1,y1) = (2,9) and m = 4 into

y - y1 = m(x - x1)

 y - 9 = 4(x - 2)

That's it in point-slope form.  

If they had said to write your answer in slope-intercept 
form you would have had to take it further to the form y=mx+b:
 y-9=4x-8
   y=4x+1

Edwin

Answer by stanbon(75887)

find the equation of the line that passes through the point (2,9) and has slope 4. write your answer in point-slope form.
----------------
Form: y = mx + b
You have y,x, and m; solve for "b":
9 = 4*2 + b
b = 1
-----------
Equation:
y = 4x+1
==============
Cheers,
Stan H.